OpenStudy (anonymous):

I have a question!! What is the value of d + e + f when d = 20, e = –4 and f = –2? A. –26 B. –22 C. 14 D. 18 I will medal

OpenStudy (ihelpyoulearn):

C!

OpenStudy (anonymous):

d+e+f=20-4-2=16-2=14

OpenStudy (anonymous):

Ah ok, thanks! can I ask another?

OpenStudy (anonymous):

go for it

OpenStudy (tkhunny):

Have you considered substituting the known values and solving for anything remaining?

OpenStudy (anonymous):

So , it's 20+-4+-2.So the first thing is to add to negative numbers.-4+-2 equals -6. Then 20 +-6 equals14 . The answer is C=14

OpenStudy (anonymous):

Ah ok thank guys!

OpenStudy (anonymous):

my computer froze can i ask second question?

OpenStudy (anonymous):

Which statement is true? A. –14 + 21 = –24 + 17 B. 16 + (–6) > 24 + (–13) C. –14 + (–10) > 2 + (–28) D. –10 + (–8) > –12 + (–4)

OpenStudy (anonymous):

yw can i get a medal for helping :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

Thats the question

OpenStudy (anonymous):

I see it

OpenStudy (anonymous):

its not A or B

OpenStudy (tkhunny):

Dear @MyPhilHillName Please show your work. Do NOT submit problems with none of your own work shown.

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

24+ -3 = 11 and 16 + -6 = 11

OpenStudy (anonymous):

im sorry I cant find the answer to that one XD

OpenStudy (anonymous):

a was, –14 + 21 which = 7

OpenStudy (anonymous):

need help with a different one?

OpenStudy (anonymous):

let me find C wait

OpenStudy (tkhunny):

You can give out more medals for good service if you post one question on one post and don't double up!

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

Ok i got it its D

OpenStudy (anonymous):

Because –10 + (–8) = -18 and –12 + (–4) = -16 and the second is smaller so >

OpenStudy (ihelpyoulearn):

@tkhunny do you ever help or just going around picking at people? haha.

OpenStudy (tkhunny):

I never pick on people. I do teach etiquette and encourage learning.